OpenStudy (anonymous):

if f(x)=x^3+ax^2+bx-3 and if f(1) =4 and f(-1)=-6 what is the value of 2a+b

OpenStudy (danjs):

you can form 2 equations involving 'a' and 'b' from those two points they give you

OpenStudy (danjs):

2 equations/2 unknowns , solve able

OpenStudy (danjs):

x = 1 , y = f(1) = 4 x=-1, y = f(-1) = -6

OpenStudy (thadyoung):

well geesh, let them try it, don't just do it all for them.

OpenStudy (danjs):

That is the first, do the same for the second, and solve for the value of a and b 4 = 1^3 + a*(1)^2 +b*(1)-3

OpenStudy (anonymous):

a=6-b?